OpenStudy (anonymous):

What is the correct way to write 0.28 as a fraction? Its not 7/25

OpenStudy (anonymous):

it should be, what else would it be?

OpenStudy (anonymous):

14/50?

OpenStudy (anonymous):

the answers are 28/100, 28/10, 2 and 8/100, or 2 and 8/10

OpenStudy (anonymous):

28/100

OpenStudy (nurali):

.28 = 28/100 = 14/50 = 7/25

OpenStudy (anonymous):

Yes it is 7/25 is correct

OpenStudy (anonymous):

whats 44.75 to the nearest tenth

OpenStudy (anonymous):

44.8 because the tenth value is the first # after the decimal ._ and in order to round it you have to see the #behind it which is 5 (.__) so your first number will increse ther value of 7--> to 8

OpenStudy (anonymous):

*increase